Find the function f+g, f-g, fg and f/g
f(x)= sqrt(x-1); g(x)=x^3+1

Answer by josgarithmetic(39621)

------------------------------------------
Find the function f+g, f-g, fg and f/g
f(x)= sqrt(x-1); g(x)=x^3+1
------------------------------------------


f+g means f(x)+g(x).

f-g means f(x)-g(x).

fg means f(x)g(x).

f/g means f(x)/g(x).

Substitute your given functions and simplify.